# Reinforcing VLMs to Use Tools for Detailed Visual Reasoning Under Resource Constraints

### v1

@@ 4 Related Work
4.1 Reasoning with VLMs
@@
As VLMs naturally inherit the reasoning capabilities from their LLM backbone, researchers attempt to teach VLMs with chain-of-thought reasoning traces to ask the model to think before answer [14, 2]. Since the breakthrough of Deepseek-r1 [19, 6], researchers have been focusing on bringing the success of r1-like GRPO training paradigms to VLMs to improve models’ visual reasoning performance. Recent research found that VLMs can achieve superior reasoning capabilities in vision-intensive mathematical tasks through GRPO [8, 34, 21, 30, 4, 26]. However, existing work often requires expensive human annotation or distillation from larger models for generalization. In the meantime, limited effort has been spent on improving VLMs’ vision-oriented reasoning capabilities, where the model can leverage task-specific details residing in images.
4.2 Fine-grained Processing of VLMs
@@
Recent research has pointed out that VLMs often neglect detailed information in images [5], and have blind faith in textual modality [18, 3]. To address this issue, inspired by humans’ behavior in recognizing information from images [32], a series of visual search-based improvements on VLMs have been proposed. Even though such methods can inject additional information of salient regions in images to the VLM, they rely on extrinsic visual search components with complex searching algorithms [29, 20, 11], thus limiting the efficiency of the system. Meanwhile, another line of research attempts to convert the input image into multi-resolution crops and feed them to the VLM end-to-end [23, 13, 35, 27]. Although these methods successfully provide the VLM with image details to tackle downstream tasks, there is computational overhead introduced by the image tokens of the cropped patches. At the same time, these solutions crop the original image generally without recognizing the importance of specific regions in the image to tackle downstream tasks, which may limit their ability to help VLMs solve real-world problems.
4.3 Multimodal Tool use with LLMs
@@
Researchers have been attempting to enable language models to do multimodal tasks by designing language model tool-use systems, i.e., developing systems that treat language models as black boxes to call APIs or use various tools [12, 16]. Without training multimodal foundational models from scratch, previous research has found that iteratively prompting LLMs with vision foundation models can successfully perform numerous vision-language tasks [31, 28, 7]. However, there is only limited research into applying the same approach to vision language models to enhance their vision-oriented reasoning capabilities [25]. A number of researchers have recently applied reinforcement learning to LLMs, including with successful tool-use learning [10, 17]. These efforts have been primarily limited to text-only models, leaving much to be explored in the multimodal domain.
The ideas and results presented in this paper were developed independently and concurrently with those of DeepEyes [36], which corroborate some of our findings. However, our focus is primarily on efficient understanding of visual detail with tool use under constrained resources.
@@ 5 Conclusion
Appendix A System and user prompts
We used the following system prompt:
You may call any of the tools exactly one time. You have access to the
following tools to help solve problems:
@@ Appendix A System and user prompts
1. Start by thinking through your reasoning inside <think> tags. Then
either return your answer inside <answer> tags, or use a tool inside
<tool> tags.
2. If needed, use a tool by writing its arguments inside <tool> tags.
Use one line for each argument in the format ’key: value’. The first
line must be ’name: <tool_name>’.
@@
Tools expect specific arguments. Follow the examples carefully for the
required keys and expected value formats.
@@
We used the following user prompt:
The image size is {image_size}. Please thoroughly think through the question and refine your answer while thinking. You should try to collect the visual evidence you need to support your answer. Then, provide your answer. The answer (which you will provide in the <answer> </answer> tags) should be a single word or phrase directly answering the question. Question: {question}

### v3

@@ 4 Related Work
Reasoning with VLMs
@@
As VLMs naturally inherit the reasoning capabilities from their LLM backbone, researchers attempt to teach VLMs with chain-of-thought reasoning traces to ask the model to think before answer (Luan et al. 2024; Bai et al. 2025). Since the breakthrough of DeepSeek-R1 (Shao et al. 2024; Guo et al. 2025), researchers have been focusing on bringing the success of R1-like GRPO training paradigms to VLMs to improve models’ visual reasoning performance. Recent research found that VLMs can achieve superior reasoning capabilities in vision-intensive mathematical tasks through GRPO (Huang et al. 2025; Zhan et al. 2025; Shen et al. 2025; Yang et al. 2025; Deng et al. 2025b; Wang et al. 2025b). However, existing work often requires expensive human annotation or distillation from larger models for generalization. In the meantime, limited effort has been spent on improving VLMs’ vision-oriented reasoning capabilities, where the model can leverage task-specific details residing in images.
Fine-grained Processing of VLMs
@@
Recent research has pointed out that VLMs often neglect detailed information in images (Gou et al. 2024), and have blind faith in textual modality (Rahmanzadehgervi et al. 2024; Deng et al. 2025a). To address this issue, inspired by humans’ behavior in recognizing information from images (Yang et al. 2020), a series of visual search-based improvements on VLMs have been proposed. Even though such methods can inject additional information of salient regions in images to the VLM, they rely on extrinsic visual search components with complex searching algorithms (Wu and Xie 2024; Shen et al. 2024; Li et al. 2025), thus limiting the efficiency of the system. Meanwhile, another line of research attempts to convert the input image into multi-resolution crops and feed them to the VLM end-to-end (Thapa et al. 2024; Liu et al. 2024; Zhang et al. 2024; Wang et al. 2025c). Although these methods successfully provide the VLM with image details to tackle downstream tasks, there is computational overhead introduced by the image tokens of the cropped patches. At the same time, these solutions crop the original image generally without recognizing the importance of specific regions in the image to tackle downstream tasks, which may limit their ability to help VLMs solve real-world problems.
Multimodal Tool use with LLMs
@@
Researchers have been attempting to enable language models to do multimodal tasks by designing language model tool-use systems, i.e., developing systems that treat language models as black boxes to call APIs or use various tools (Li et al. 2023; Peng et al. 2023). Without training multimodal foundational models from scratch, previous research has found that iteratively prompting LLMs with vision foundation models can successfully perform numerous vision-language tasks (Yang et al. 2023; Wu et al. 2023; Gupta and Kembhavi 2023). However, there is only limited research into applying the same approach to vision language models to enhance their vision-oriented reasoning capabilities (Wang et al. 2025a). A number of researchers have recently applied reinforcement learning to LLMs, including with successful tool-use learning (Jin et al. 2025; Qian et al. 2025). These efforts have been primarily limited to text-only models, leaving much to be explored in the multimodal domain.
More recently, DeepEyes (Zheng et al. 2025), corroborates some of our findings on using RL to encourage tool usage for vision. However, our work is distinct in its focus on identifying practical approaches and best practices for tool use under constrained computational resources, targeting smaller models and efficient training strategies.
@@ 5 Conclusion
Acknowledgments
This research was funded by Groundlight AI.
Appendix A Prompts
Below, you can find:
  1. 1.
    The system prompt, which describes the desired response format and how tools can be invoked.
  2. 2.
    The zoom tool description, which provides the model with the appropriate context to use the tool.
  3. 3.
    The user prompt, which shows how we prompted the model to answer each question during training and evaluation.
System Prompt You may call any of the tools exactly one time. You have access to the following tools to help solve problems:
@@ Appendix A Prompts
1. Start by thinking through your reasoning inside <think> tags. Then either return your answer inside <answer> tags, or use a tool inside <tool> tags.
2. If needed, use a tool by writing its arguments inside <tool> tags. Use one line for each argument in the format "key: value". The first line must be "name: <tool_name>".
@@
Tools expect specific arguments. Follow the examples carefully for the required keys and expected value formats.
@@
Zoom Tool Description Zoom Tool: Returns the image zoomed in on the specified keypoint.
This is useful to see a region of interest with higher clarity, like for reading text or identifying objects.
Args:
   keypoint: list[int], the [x, y] coordinates of the point to center the zoom on.
   Must be within the image boundaries.
Returns:
   The image zoomed in on the specified keypoint.
Examples:
   <tool>
   name: zoom
   keypoint: [500, 400]
   </tool>
<tool>
   name: zoom
   keypoint: [23, 44]
   </tool>
<tool>
   name: zoom
   keypoint: [723, 461]
   </tool>
User Prompt The image size is {image_size}.
Please thoroughly think through the question and refine your answer while thinking. You should try to collect the visual evidence you need to support your answer. Then, provide your answer. The answer (which you will provide in the <answer> </answer> tags) should be a single word or phrase directly answering the question.
Question: {question}